Circular No. 6154

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA AQUILAE 1995 C. G. Mason and R. D. Gehrz, University of Minnesota; and C. E.

Woodward, University of Wyoming, report magnitudes for N Aql 1995 obtained on Mar. 23.6 UT with a bolometer at the Wyoming Infrared Observatory’s 2.3-m telescope, using a 10′′ beam and a 38′′ throw: J = 8.00 ± 0.13, H = 7.21 ± 0.03, K = 5.17 ± 0.02, L = 2.79 ± 0.04, M = 1.76 ± 0.05, N = 0.75 ± 0.20, [7.8 μm] = 0.84 ± 0.06, [8.7 μm] = −0.26 ± 0.09, [9.8 μm] = 0.90 ± 0.19, [10.3 μm] = 1.58 ± 0.15. They add: “The measurements conform to a 970-K blackbody, suggesting recent dust formation. The temperature is close to the 1000- to 1200-K dust condensation temperatures observed in other novae. We encourage visual observers to monitor this nova, as it may be expected that the dust will cause a DQ Her-type transition in the light curve.”

RX J0558+53 A. Allan, K. Horne, and R. W. Hilditch, St. Andrews University, write:

“Based on V - and R-band CCD light curves between 1994 Nov. 24 and 1995 Feb. 22, we find a period of 272.7277(4) s for RX J0558+53 (cf. IAUC 6129). Power spectra show a peak at twice this period, suggesting that the true spin period is 545.4555(8) s.”

NOVA CENTAURI 1995 F. Jablonski, Laboratorio Nacional de Astrofisica (LNA), Itajuba, re-

ports: “UBV(RI)_{c} photometry, obtained with the CNPq/LNA 0.6-m Zeiss telescope, yields the following average photometry for N Cen 1995: Mar. 16.27 UT, U −B = −0.55, B−V = +0.33, V = 9.89, V −R = +1.36, V −I = +1.48; 17.19, −0.58, +0.29, 9.75, +1.34, +1.45; 18.21, −0.56, +0.36, 9.97, +1.36, +1.43; 19.19, –, –, 10.04, –, –. The photometric errors are 0.01 mag in all bands, except U − B (0.02 mag); photometric transformations subject to the constraints discussed by Jablonski et al. (1994, PASP 106, 1172). CCD spectra (resolution ∼ 20000) obtained on Mar. 18.1 with the coudé spectrograph at the 1.6-m telescope show a strong interstellar feature at 589.0/589.6 nm. The total equivalent widths of the components of the Na I doublet are 0.097 and 0.0742 nm. Each feature is resolved into two components with residual velocities with respect to the local standard of rest of −22 and −1 km/s, respectively.”

1995 April 3 Daniel W. E. Green